OpenStudy (anonymous):

Convert rectangular coordinates (0,-sqrt(3)) to polar coordinates. Use radians and choose the angle between the interval (-pi, pi)

OpenStudy (anonymous):

\(r=\sqrt{3} \ \theta=\pi\)

OpenStudy (anonymous):

It says that theta doesn't equal pi

OpenStudy (anonymous):

My fault: \(\theta =-\pi/2\)